# Low dimensional representation of multi-patient flow cytometry datasets using optimal transport for measurable residual disease detection in leukemia

### v2

@@ Abstract
Representing and quantifying Measurable Residual Disease (MRD) in Acute Myeloid Leukemia (AML), a type of cancer that affects the blood and bone marrow, is essential in the prognosis and follow-up of AML patients. As traditional cytological analysis cannot detect leukemia cells below 5%, the analysis of flow cytometry dataset is expected to provide more reliable results. In this paper, we explore statistical learning methods based on optimal transport (OT) to achieve a relevant low-dimensional representation of multi-patient flow cytometry measurements (FCM) datasets considered as high-dimensional probability distributions. Using the framework of OT, we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. After this quantization step, the visualization of the intra and inter-patients FCM variability is carried out by embedding low-dimensional quantized probability measures into a linear space using either Wasserstein Principal Component Analysis (PCA) through linearized OT or log-ratio PCA of compositional data. Using a publicly available FCM dataset and a FCM dataset from Bordeaux University Hospital, we demonstrate the benefits of our approach over the popular kernel mean embedding technique for statistical learning from multiple high-dimensional probability distributions. We also highlight the usefulness of our methodology for low-dimensional projection and clustering patient measurements according to their level of MRD in AML from FCM. In particular, our OT-based approach allows a relevant and informative two-dimensional representation of the results of the FlowSom algorithm, a state-of-the-art method for the detection of MRD in AML using multi-patient FCM.
@@ 1 Introduction
Acute Myeloid Leukemia (AML) is a cancer of the blood and bone marrow when the development of blood stem cells (immature cells) into mature cells is altered, and the number of these immature cells, called myeloid blasts, increases. The diagnosis of AML consists in the detection of blasts which invade the bone marrow to great amounts, and it is typically assessed with the cytological analysis of the patient’s bone marrow. The goal of the chemotherapy treatment is therefore to eradicate the blasts population: 80% of patients reach complete remission after one cycle of chemotherapy, however 50% relapse [40]. The definition of complete remission states that the bone marrow should contain less than 5% of blasts. However, detecting smaller numbers of persisting residual leukemic cells, called Measurable Residual Disease (MRD), is critical for disease monitoring and relapse prognosis. This motivates the need to catch the presence of blasts at very low thresholds (between 0.1% and 0.001%). The detection and the quantification of MRD in AML has been widely studied as a predictive and prognostic tool [18, 23, 37, 38, 16, 24]. Molecular biology and flow cytometry are the main two techniques currently available to achieve informative sensitivity levels for anticipating the disease’s progression. While molecular biology techniques for the assessment of MRD were the first to be developed and have a recognized decision-making value [34, 5], it has limited applicability (only available for 40% of the patients) due to the lack of sufficiently specific molecular targets in 60% of patients.
@@
Flow cytometry is a technology which allows the fast analysis of multiple parameters of a large population of cells from a biological sample. The interaction of the cells with the light emitted from different lasers enables the evaluation of two types of characteristics: the structure of the cell (size, granularity) and the type of proteins expressed on the surface of the analyzed cell. For a given sample, each parameter is measured for each cell. One of the main goal of flow cytometry data analysis is then to identify the different types of cells in a biological sample, and to be able to locate regions of interest. Although flow cytometry dataset is widely used in practice, analysis of the datasets is not straightforward, as the large amount of data they generate and the dimensionality of the observations pose significant scientific challenges for their interpretation for MRD assessment purposes [40]. Additionally surface markers used to characterize AMLs have highly variable expression levels from patient to patient, ranging from overexpression to expression negativity, and it is the combination of these abnormalities that defines the leukemic phenotype.
@@
The current preferred method for identifying cells is the so-called manual gating. It is based on the successive projection of the point cloud of the cytometry measurements according to two selected parameters. The experts then partition the space according to the two selected parameters, visually identifying regions of high density. In the diagnosis of AML, the process of manual gating is well-suited and largely employed to the detection of blasts (tumor cells) which invade the body to great amounts (at thresholds of $10^{11}$ to $10^{12}$), allowing easy identification of the tumor cell subpopulation. However, it presents a few shortcomings: it is highly dependent on the biologist’s expertise and it is extremely time-consuming. Moreover, the significant variability in flow cytometry data is a limiting factor. This variability is induced both by the biological heterogeneity between the samples and by technical details such as the cytometer performance and the data acquisition parameters. Finally, because of the very low thresholds of MRD, the identification based on manual gating of cellular subpopulation of blasts in flow cytometry data during follow-up of patients treated with chemotherapy is made tricky.
@@
For a given patient, a single dataset of flow cytometry is usually made of measurements using 10 to 20 parameters from hundreds of thousands of cells. This motivates the need to develop dimensionality reduction tools for representing variations of interest within datasets acquired from an individual patient and across a population of patients using multiple flow cytometry measurements (FCM). To overcome the shortcomings of a manual approach to data gating in FCM, the development of automatic inference methods is currently a very active field of research in computational statistics and bioinformatics. Several methods have been introduced to automate the gating of cytometry data [43]. One can cite unsupervised algorithms such as FlowMeans [1], based on K-means, Cytometree [17], based on binary trees, or [22] based on Dirichlet process mixtures. There also exist supervised algorithms such as DeepCyTOF [25] or flowlearn [26]. Regarding MRD detection for AML, [40] proposes a method using FlowSOM [39], a visualization and clustering method based on self-organising maps. In this paper, we will denote MRD-FlowS the MRD measure calculated by this method. On the other hand, MRD-BioM will designate the MRD value assessed with molecular biology. In [31], the authors also use FlowSOM to detect MRD for chronic lymphocytic leukemia.
@@
Optimal Transport (OT) provides a versatile and powerful framework for addressing a diverse array of problems involving point cloud modeled as discrete probability measures. In the last decade, the use of computational OT [33] in data science and the related notion of Wasserstein distance have gained an increasing interest in statistics (see e.g. [7, 32] for recent overviews), in particular for dimension reduction of a set of probability measures using the notion of Wasserstein Principal Component Analysis (PCA) [8, 13]. OT has also proved useful for understanding various phenomena in biology [6, 12, 20, 36]. For the purpose of automatic gating from flow cytometry data, OT has been previously leveraged for quantifying the distance between two point clouds, taking into account the geometry of the data in high-dimensional spaces, giving rise to two OT-based algorithms: optimalFlow [19] proposes a method for supervised gating, and CytOpt [21] estimates the different cell proportions of a population from cytometry measurements. More recently, [27] introduced a method based on linear optimal transport for the classification of point clouds and showed its applications to flow cytometry datasets from AML patients.
@@ 1.1 Contributions
We explore a statistical methodology based on computational OT for embedding into a low-dimensional linear space a set of flow cytometry measurements (FCM) modeled as high-dimensional probability distributions. Our approach is suitable for the study of datasets that consist of multiple FCM from various patients with possibly several measurements for a given patient (e.g.  acquired at different time or from different cytometers). Using the notion of Wasserstein barycenter [2], we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. This quantization step allows to reduce the size of the support of the observed data viewed as high-dimensional discrete probability measures. The resulting low-dimensional support quantized probability measures are then embedded into a linear space using either Wasserstein PCA [8, 13] through linearized OT [11, 42, 29, 27] or log-ratio PCA of compositional data [3]. This enables the analysis of intra and inter-patients FCM variability. Using a publicly available dataset and a dataset provided by Bordeaux University Hospital, this approach is first shown to achieve a meaningful clustering of patients according to their FCM. We also demonstrate the benefits of our methodology over the popular kernel mean embedding (KME) technique [30] for statistical learning from multiple high-dimensional probability distributions. Secondly, we study applications to the assessment of MRD in AML. In particular, our OT-based approach is found to provide a relevant and informative two-dimensional representation of the results obtained with the FlowSom algorithm (MRD-FlowS), a robust method with growing adoption in the field for the detection of MRD in AML using FCM. This representation could be a valuable tool to help biologists confirm or infirm the relapse prognosis of AML patients.
@@ 1.2 Outline of the paper
1.3 Availability and implementation
All the analysis carried out in this work to analyze multiple FCM are reproducible through the use of Python codes that are available at https://github.com/erellgachon/CytoLOT.
@@ 2.1 DATAML-Bordeaux study
From this study, we have kept 182 patients and 22 normal bone marrow (NBM). For each patient, two or three FCM are available: one diagnosis ($\sim$50000 cells) and one or two follow-ups (up to 500000 cells). In total, we have 512 FCM. For each follow-up, we dispose of its associated MRD-FlowS estimation with the FlowSOM method. The MRD-BioM measure is also given when available (108 out of 314 follow-ups), which we consider as ground truth. There are 11 markers: 2 are structural (FS INT, SS INT) and 9 are protein expressions (CD34, CD13, CD38, CD7, CD33, CD56, CD117, HLA-DR, CD45). Hence, the datasets consist of multiple point clouds in dimension $d=11$. We represent the $N=512$ FCM as high-dimensional discrete probability measures to leverage OT statistical learning tools. For $1\leq i\leq N$, we represent the $i$-th dataset $X^{i}=(X^{i}_{1},\cdots,X^{i}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}}$ with $m_{i}$ the number of cells in this dataset, as a discrete measure $\mu^{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\delta_{X_{j}^{i}}$.
The cytometry files are stored in both FCS and CSV format and have already undergone standard pre-processing, i.e. transformation, compensation and normalization.
@@ 2.2 HIPC dataset
Using a second dataset, that is publicly available on ImmuneSpace [10], we illustrate the benefits of our methodology to cluster FCM data by analyzing observations from the T-cell panel of the Human Immunology Project Consortium (HIPC). Seven laboratories stained three replicates (denoted A, B, and C) of three cryo-preserved biological samples denoted patient 1, 2, and 3 (e.g. cytometry measurements from the Stanford laboratory for replicate C from patient 1). Each resulting dataset, that consists of the measurements of 7 markers for approximately $100000$ cells, thus leads to a set of FCM consisting of $N=7\times 3\times 3=63$ point clouds in dimension $d=7$. For this dataset, a relevant clustering consists of three groups, each comprising data collected from a single patient.
@@ 2.3 The FlowSOM method
FlowSOM [39] is a visualization and clustering method that is specially designed for analyzing flow cytometry data with Self-Organizing Maps (SOMs). This algorithm has been leveraged in [40] for the evaluation of MRD in AML.
@@
The method in [40] to evaluate MRD in a follow-up of a patient undergoing treatment first consists in merging three cytometry entry files: a normal bone marrow (from an arbitrary, healthy donor), the diagnosis of the patient with AML, and his follow-up. The authors in [40] then perform a FlowSOM on the merged files, the last step of which results in an MST. They visualize the weights on the MST for each entry file. By comparing the three MST globally, they are able to identify the Nodes of Interests (NoIs) that are decisive in assessing the MRD present in the measurements. It is reported in [40] that this method shows 80% concordance with MRD-Bio, that is with the ground truth results. More precisely, MRD-FlowS yields 85% specificity and 69% sensitivity compared to MRD-BioM.
@@ 2.4.2 Dimension reduction via mean measure quantization
The Wasserstein barycenter problem with varying weights (5) is equivalent to the $K$-means quantization of the mean measure $\overline{\mu}=\frac{1}{N}\sum_{i=1}^{N}\mu^{i}$, that is,
@@ 2.4.3 An embedding with the OT linearization
The Wasserstein space consisting of probability measures admitting a moment of order 2 endowed with the 2-Wasserstein distance has nice geometric properties such as existence of geodesics and a pseudo-Riemannian structure [4]. However, it is not a Hilbert space [33], and tools such as PCA or LDA for statistical learning cannot be directly applied on raw probability measures. The OT linearization then consists in embedding probability measures in a linear space, allowing to apply standard statistical tools [11, 42]. Given a set of probability measures $(\mu_{i})_{1\leq i\leq N}$ and an absolutely continuous reference measure $\mu$, linearized OT consists in embedding the $\mu_{i}$’s in the tangent space $\mathcal{T}_{{\mu}}$ at $\mu$, defined as a subspace of the Hilbert space $L^{2}({\mu},\mathbb{R}^{d})=\{v:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}~{}|~{}\int_{\mathbb{R}^{d}}\|v\|^{2}\mathrm{d}{\mu}<\infty\}$, endowed with the weighted $L^{2}$ inner product $\langle T_{1},T_{2}\rangle_{L^{2}({\mu})}=\int_{\mathbb{R}^{d}}T_{1}(x)T_{2}(x)\mathrm{d}{\mu}(x)$. A measure $\mu^{i}$ is therefore embedded into $L^{2}({\mu},\mathbb{R}^{d})$ through the following logarithm map, by analogy with the Riemannian setting :
@@
where $\nu_{\text{bar}}=\sum_{k=1}^{K}\bar{a}_{k}\delta_{x_{k}}$ and $\nu^{i}=\sum_{k=1}^{K}a_{k}^{i}\delta_{x_{k}}$. Note that the barycentric mappings $(T^{i})_{i=1}^{N}$ are well defined on the support of the reference measure $\nu_{\text{bar}}$.

### v3

@@ Abstract
Representing and quantifying Measurable Residual Disease (MRD) in Acute Myeloid Leukemia (AML), a type of cancer that affects the blood and bone marrow, is essential in the prognosis and follow-up of AML patients. As traditional cytological analysis cannot detect leukemia cells below 5%, the analysis of flow cytometry datasets is expected to provide more reliable results. In this paper, we explore statistical learning methods based on optimal transport (OT) to achieve a relevant low-dimensional representation of multi-patient flow cytometry measurements (FCM) datasets considered as high-dimensional probability distributions. Using the framework of OT, we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. After this quantization step, the visualization of the intra- and inter-patient FCM variability is carried out by embedding low-dimensional quantized probability measures into a linear space using either Wasserstein Principal Component Analysis (PCA) through linearized OT or log-ratio PCA of compositional data. Using a publicly available FCM dataset and a FCM dataset from Bordeaux University Hospital, we demonstrate the benefits of our approach over the popular kernel mean embedding technique for statistical learning from multiple high-dimensional probability distributions. We also highlight the usefulness of our methodology for low-dimensional projection and clustering patient measurements according to their level of MRD in AML from FCM. In particular, our OT-based approach allows a relevant and informative two-dimensional representation of the results of the FlowSom algorithm, a state-of-the-art method for the detection of MRD in AML using multi-patient FCM.
Data Availability Statement: All the analysis carried out in this work to analyze multiple FCM are reproducible through the use of Python codes that are available at https://github.com/erellgachon/CytoLOT.
Correspondance: Erell Gachon1. Email : erell.gachon@math.u-bordeaux.fr.
@@ 1 Introduction
Acute Myeloid Leukemia (AML) is a cancer of the blood and bone marrow when the development of blood stem cells (immature cells) into mature cells is altered, and the number of these immature cells, called myeloid blasts, increases. The diagnosis of AML consists in the detection of blasts which invade the bone marrow to great amounts, and it is typically assessed with the cytological analysis of the patient’s bone marrow. The goal of the chemotherapy treatment is therefore to eradicate the blasts population: 80% of patients reach complete remission after one cycle of chemotherapy, however 50% relapse [43]. The definition of complete remission states that the bone marrow should contain less than 5% of blasts. However, detecting smaller numbers of persisting residual leukemic cells, called Measurable Residual Disease (MRD), is critical for disease monitoring and relapse prognosis. This motivates the need to catch the presence of blasts at very low thresholds (between 0.1% and 0.004%). The detection and the quantification of MRD in AML has been widely studied as a predictive and prognostic tool [18, 23, 40, 41, 16, 24]. Molecular biology and flow cytometry are the main two techniques currently available to achieve informative sensitivity levels for anticipating the disease’s progression. While molecular biology techniques for the assessment of MRD were the first to be developed and have a recognized decision-making value [36, 5], they have limited applicability (only available for 40% of the patients) due to the lack of sufficiently specific molecular targets in 60% of patients.
@@
Flow cytometry is a technology which allows the fast analysis of multiple parameters of a large population of cells from a biological sample. The interaction of the cells with the light emitted from different lasers enables the evaluation of two types of characteristics: the structure of the cell (size, granularity) and the type of proteins expressed on the surface of the analyzed cell. For a given sample, each parameter is measured for each cell. One of the main goals of flow cytometry data analysis is then to identify the different types of cells in a biological sample, and to be able to locate regions of interest. Although flow cytometry datasets are widely used in practice, analysis of the datasets is not straightforward, as the large amount of data they generate and the dimensionality of the observations pose significant scientific challenges for their interpretation for MRD assessment purposes [43]. Additionally surface markers used to characterize AMLs have highly variable expression levels from patient to patient, ranging from overexpression to expression negativity, and it is the combination of these abnormalities that defines the leukemic phenotype.
@@
The current preferred method for identifying cells is the so-called manual gating. It is based on the successive projection of the point cloud of the cytometry measurements according to two selected parameters. The experts then partition the space according to the two selected parameters, visually identifying regions of high density. In the diagnosis of AML, the process of manual gating is well-suited and largely employed to the detection of blasts (tumor cells) which invade the body to great amounts (at thresholds of $10^{11}$ to $10^{12}$ cells in the body), allowing easy identification of the tumor cell subpopulation. However, it presents a few shortcomings: it is highly dependent on the biologist’s expertise and it is extremely time-consuming. Moreover, the significant variability in flow cytometry data is a limiting factor. This variability is induced both by the biological heterogeneity between the samples and by technical details such as the cytometer performance and the data acquisition parameters. Finally, because of the very low thresholds of MRD, the identification based on manual gating of cellular subpopulation of blasts in flow cytometry data during follow-up of patients treated with chemotherapy is made tricky.
@@
For a given patient, a single dataset of flow cytometry is usually made of measurements using 10-20 parameters from hundreds of thousands of cells. This motivates the need to develop dimensionality reduction tools for representing variations of interest within datasets acquired from an individual patient and across a population of patients using multiple flow cytometry measurements (FCM). To overcome the shortcomings of a manual approach to data gating in FCM, the development of automatic inference methods is currently a very active field of research in computational statistics and bioinformatics. Several methods have been introduced to automate the gating of cytometry data [46]. One can cite unsupervised algorithms such as FlowMeans [1], based on K-means, Cytometree [17], based on binary trees, or [22] based on Dirichlet process mixtures. There also exist supervised algorithms such as DeepCyTOF [27] or flowlearn [28]. Regarding MRD detection for AML, [43] proposes a method using FlowSOM [42], a visualization and clustering method based on self-organizing maps. In this paper, we will denote MRD-FlowS the MRD measure calculated by this method. On the other hand, MRD-BioM will designate the MRD value assessed with molecular biology. In [33], the authors also use FlowSOM to detect MRD for chronic lymphocytic leukemia.
@@
Optimal Transport (OT) provides a versatile and powerful framework for addressing a diverse array of problems involving point clouds modeled as discrete probability measures. In the last decade, the use of computational OT [35] in data science and the related notion of Wasserstein distance have gained increasing interest in statistics (see e.g. [8, 34] for recent overviews), in particular for dimension reduction of a set of probability measures using the notion of Wasserstein Principal Component Analysis (PCA) [9, 13]. OT has also proved useful for understanding various phenomena in biology [7, 12, 20, 39]. For the purpose of automatic gating from flow cytometry data, OT has been previously leveraged for quantifying the distance between two point clouds, taking into account the geometry of the data in high-dimensional spaces, giving rise to two OT-based algorithms: optimalFlow [19] proposes a method for supervised gating, and CytOpt [21] estimates the different cell proportions of a population from cytometry measurements. More recently, [29] introduced a method based on linear optimal transport for the classification of point clouds and showed its applications to flow cytometry datasets from AML patients.
@@ 1.1 Contributions
The novelty of our work compared to previous studies [6, 45] revolving around linear optimal transport and PCA is a statistical methodology based on computational OT for embedding into a low-dimensional linear space a set of flow cytometry measurements (FCM) modeled as high-dimensional probability distributions. Our approach is suitable for the study of datasets that consist of multiple FCM from various patients with possibly several measurements for a given patient (e.g., acquired at different time or from different cytometers). Using the notion of Wasserstein barycenter [2], we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. This quantization step allows to reduce the size of the support of the observed data viewed as high-dimensional discrete probability measures. The resulting low-dimensional support quantized probability measures are then embedded into a linear space using either Wasserstein PCA [9, 13] through linearized OT [45, 6, 31, 29] or log-ratio PCA of compositional data [3]. This enables the analysis of intra- and inter-patient FCM variability. Using a publicly available dataset and a dataset provided by Bordeaux University Hospital, this approach is first shown to achieve a meaningful clustering of patients according to their FCM. We also demonstrate the benefits of our methodology over the popular kernel mean embedding (KME) technique [32] for statistical learning from multiple high-dimensional probability distributions. Secondly, we study applications to the assessment of MRD in AML. In particular, our OT-based approach is found to provide a relevant and informative two-dimensional representation of the results obtained with the FlowSom algorithm (MRD-FlowS), a robust method with growing adoption in the field for the detection of MRD in AML using FCM. This representation could be a valuable tool to help biologists confirm or invalidate the relapse prognosis of AML patients.
@@ 2.1 DATAML-Bordeaux Study
From this study, we have kept 182 patients and 22 normal bone marrow (NBM). For each patient, two or three FCM are available: one diagnosis ($\sim$50,000 cells) and one or two follow-ups (up to 500,000 cells). In total, we have 512 FCM. For each follow-up, we dispose of its associated MRD-FlowS estimation with the FlowSOM method.
For the follow-ups, the MRD-BioM measure is also given when available (108 out of 314 follow-ups), which we consider as ground truth. Concerning molecular biology, RNAs were purified from Bone Marrow mononuclear cells using Trizol reagent (Invitrogen-Thermo Fischer, Carlsbad, CA, USA). Reverse transcription was performed using the SuperScript IV VILO Master Mix (Lifetechnologies-Thermo Fischer, Carlsbad, CA, USA). The assessment of transcript levels was performed on a Light Cycler 480 (Roche LifeScience, Penzberg, Germany) with a specific RQ-PCR assay as previously described in [26]. MRD levels were reported as the normalized values of fusion transcripts or NPM1mut copy number/ABL copy number $\times$ 100 (%). The quantitative detection limit of the assays was 0.02% for NPM1 mutations. The achievement of MRD levels below this threshold was defined as a negative MRD.
There are 11 markers in the flow cytometry datasets : 2 are structural (FS INT, SS INT) and 9 are protein expressions (CD34, CD13, CD38, CD7, CD33, CD56, CD117, HLA-DR, CD45). Hence, the datasets consist of multiple point clouds in dimension $d=11$. We represent the $N=512$ FCM as high-dimensional discrete probability measures to leverage OT statistical learning tools. For $1\leq i\leq N$, we represent the $i$-th dataset $X^{i}=(X^{i}_{1},\cdots,X^{i}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}}$ with $m_{i}$ the number of cells in this dataset, as a discrete measure $\mu^{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\delta_{X_{j}^{i}}$.
The cytometry files are stored in both FCS and CSV format and have already undergone standard pre-processing, i.e., transformation, compensation and normalization.
@@ 2.2 HIPC dataset
Using a second dataset, that is publicly available on ImmuneSpace [11], we illustrate the benefits of our methodology to cluster FCM data by analyzing observations from the T-cell panel of the Human Immunology Project Consortium (HIPC). Seven laboratories stained three replicates (denoted A, B, and C) of three cryo-preserved biological samples denoted patient 1, 2, and 3 (e.g., cytometry measurements from the Stanford laboratory for replicate C from patient 1). Each resulting dataset, that consists of the measurements of 7 markers for approximately $100000$ cells, thus leads to a set of FCM consisting of $N=7\times 3\times 3=63$ point clouds in dimension $d=7$. The markers are the following : CCR7, CD4, CD45RA, CD3, HLADR, CD38, CD8. For this dataset, a relevant clustering consists of three groups, each comprising data collected from a single patient.
@@ 2.3 The FlowSOM Method
FlowSOM [42] is a visualization and clustering method that is specifically designed for analyzing flow cytometry data with Self-Organizing Maps (SOMs). This algorithm has been leveraged in [43] for the evaluation of MRD in AML.
@@
The method in [43] to evaluate MRD in a follow-up of a patient undergoing treatment first consists of merging three cytometry entry files: a normal bone marrow (from an arbitrary, healthy donor), the diagnosis of the patient with AML, and his follow-up. The authors in [43] then perform a FlowSOM on the merged files, the last step of which results in an MST. They visualize the weights on the MST for each entry file. By comparing the three MSTs globally, they are able to identify the Nodes of Interest (NoIs) that are decisive in assessing the MRD present in the measurements. It is reported in [43] that this method shows 80% concordance with MRD-BioM, that is with the ground truth results. More precisely, MRD-FlowS yields 85% specificity and 69% sensitivity compared to MRD-BioM.
@@ 2.4.1 Optimal Transport and Wasserstein Barycenter
However, using OT in problems involving point clouds suffers from high computational costs as it has complexity $O(n^{3}\log(n))$ [35] for points clouds of $n$ points. Hence, when the number of points per clouds is larger than $10^{5}$, solving the exact OT problem becomes too expensive.
@@ 2.4.2 Dimension Reduction via Mean Measure Quantization
When the $\mu^{i}$’s are absolutely continuous w.r.t. the Lebesgue measure, the Wasserstein barycenter problem with varying weights (5) is equivalent to the $K$-means quantization of the mean measure $\overline{\mu}=\frac{1}{N}\sum_{i=1}^{N}\mu^{i}$, that is,
@@ 2.4.3 An Embedding with the OT Linearization
The Wasserstein space consisting of probability measures admitting a moment of order 2 endowed with the 2-Wasserstein distance has nice geometric properties such as existence of geodesics and a pseudo-Riemannian structure [4]. However, it is not a Hilbert space [35], and tools such as PCA or LDA (Linear Discriminant Analysis) for statistical learning cannot be directly applied on raw probability measures. The OT linearization [45] then consists in embedding probability measures in a linear space, allowing to apply standard statistical tools. In particular it has been used to perform PCA on images modeled as probability measures [6, 45]. Given a set of probability measures $(\mu_{i})_{1\leq i\leq N}$ and an absolutely continuous reference measure $\mu$, linearized OT consists in embedding the $\mu_{i}$’s in the tangent space $\mathcal{T}_{{\mu}}$ at $\mu$, defined as a subspace of the Hilbert space $L^{2}({\mu},\mathbb{R}^{d})=\{v:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}~{}|~{}\int_{\mathbb{R}^{d}}\|v\|^{2}\mathrm{d}{\mu}<\infty\}$, endowed with the weighted $L^{2}$ inner product $\langle T_{1},T_{2}\rangle_{L^{2}({\mu})}=\int_{\mathbb{R}^{d}}T_{1}(x)T_{2}(x)\mathrm{d}{\mu}(x)$. A measure $\mu^{i}$ is therefore embedded into $L^{2}({\mu},\mathbb{R}^{d})$ through the following logarithm map, by analogy with the Riemannian setting :
@@
where $\nu_{\text{bar}}=\sum_{k=1}^{K}\bar{a}_{k}\delta_{x_{k}}$ and $\nu^{i}=\sum_{k=1}^{K}a_{k}^{i}\delta_{x_{k}}$. Note that the barycentric mappings $(T^{i})_{i=1}^{N}$ are well defined on the support of the reference measure $\nu_{\text{bar}}$. Applying linearized OT to a set of $N$ (typically in the DataML-Bordeaux study, we have $N\sim 500$) FCMs would boil down to solving $N$ times the OT problem with sets of $n\sim 10^{5}$ points. This motivates the need for the previously described mean measure quantization step, which allows to reduce the size of the point clouds, and consequently decrease the computational times.